If a standard number cube or die is rolled what is the probability that a 4 does not land face up?
If a standard number cube or die is rolled, the probability that a 4 does not land face up is five out of six, or (six minus one) out of six.
Two unbiased dice are thrown Find the probability that neither double nor a total of 10 will appear?
The sample space is 62 or 36. The easier way to do this problem is to find the probability of doubles or 10 and subtract from 1. The probability of a double is 6/36 (1,1; 2,2; 3,3; 4,4; 5,5; 6,6). There are 3 ways to get a 10; 4,6; 6,4; & 5,5. We already counted 5,5 in the double so there are 2/36 we add to the 6/36 to obtain 8/36. So, 1-8/36 = 28/36 or 7/9 or 0.7778 probability that neither double nor a total of 10 will appear.
14/27 ~ 0.519 ~ 51.9%
----------------------------------------------------------------------------------------------------
EXPLANATION
For bag 1: Probability of drawing a black ball; P(B1) =
6/9. A white one; P(W1) =3/9
For bag 2: P(B2) =
5/9. P(W2) =
4/9.
Probability of drawing a black ball from each bag: P(B1UB2) =
(6/9)∙(5/9) =
10/27
Probability of drawing a white ball from each bag: P(W1UW2) =
(3/9)∙(4/9) =
4/27
The probability of drawing 2 balls of the same color is the sum of the above
probabilities:
P(2 same color balls) =
P(B1UB2) + P(W1UW2) =
10/27 +4/27 =
14/27 =
0.5185...
P(2 same color balls) ~ 0.519 ~ 51.9%
If you have a 1 percent chance of doing something, and you have 30 tries, your chance of succeeding is 26 percent, or 1 - (1 - .01)30.
A single card is drawn from a deck Find the probability of selecting a 10 or a diamond?
for the diamonds - you have a 1in4 chance (25%) 52 cards in the deck, 13 are diamonds or 13:52=1:4 for the 10s - you have 1in13 chance (7.7%) 52 cards in the deck, 4 are 10s or 4:52=1:13
What are the chances of rolling double 6 3 times in a row?
The chance of rolling a double 6 is 1/6x1/6
The chance of getting the 3 times in a row is
(1/6x1/6)3 or (1/36)3 which is 1/46,656, a very small number!
0.97. Just take it away from 1
How do you refill ace liner 502?
This isn't filled like your normal standard stapler. First, pull up the top half of the stapler (the actual part that holds the staples and basically does all the work. Second, push in the two side buttons that are about 1/4 of the way down from the front of the stapling head. This will open the staplereservoir. There is a rod inside that has a round disc that comes out the back. Grab ahold of the disc and push in then lift up. This will release the rod. Pull the rod all the way back, as far as it will go. It doesn't come out, it will stop once it reaches the farthest back position. Lift the rod up to approximately a 90 degree angle. The staples go inside the opening that is now exposed. The basically rest along a rail. Push them in as far as they go and then return the rod you just pulled out and lifted up to it's natural parallel position. Push the rod back in, allowing the spring on the rod to create the tension needed to hold the staples in place and snap the catch clip near the back of the rod (just in front of the disc) back down into place. That's it! It sounds more complicated than it is but it's really not at all once you do it one time.
Probability of picking one red card of a deck of 52 playing cards?
The probability of picking one red card of a deck of 52 playing cards is 26 out of 52, or 1 out of 2.
What is the probability of rolling a 7 or 11 with two dice?
The probability of rolling a 7 with 2 dice is 6/36; probability of rolling an 11 is 2/36. Add the two together to find probability of rolling a 7 or 11 which is 8/36 or 2/9.
Which normal distribution is also the standard normal curve?
The normal distribution would be a standard normal distribution if it had a mean of 0 and standard deviation of 1.
What are two events that cannot occur at the same time called?
Two events that cannot occur at the same time are called mutually exclusive.
If you roll a fair six-sided die 1000 times the die would come up even about 500 times. It will not be exactly 500 times, due to random probability, but it will be close. The more times you roll it, the ratio of even to odd will come closer and closer to the theoretical probability of 0.5.
The word mathematics has 11 letters; 2 are m, a, t. The number of distinguishable permutations is 11!/(2!2!2!) = 39916800/8 = 4989600.
Is the Z-test the same as Z-scores?
No, the Z-test is not the same as a Z-score. The Z-test is where you take the Z-score and compare it to a critical value to determine if the null hypothesis will be rejected or fail to be rejected.
If two events are mutually exclusive then they must be dependent?
No, if two events are mutually exclusive, they cannot both occur. If one occurs, it means the second can not occur.
The probability that 2 people have the same number is 2 out of 10
A spinner a-f what is the probability of the spinner landing on b then landing on a?
Assuming that it is a regular shaped spinner, the probability is 1/6*1/6 = 1/36
Hasty generalizing
If you toss a fair coin 250 times about how many times should it land on tails?
If you toss a fair coin 250 times , about how many times should it land on tails?